OpenStudy (anonymous):

help Solve for x. x/3-(x-1)/2>=1

OpenStudy (anonymous):

is this the complete question?

OpenStudy (anonymous):

>=

OpenStudy (anonymous):

>= what ?

OpenStudy (anonymous):

the > is over the =

OpenStudy (junyang96):

- what? lol

OpenStudy (anonymous):

yes but I need something on the right hand side after the symbol or I can't solve it

OpenStudy (anonymous):

ther is nothing on that side

OpenStudy (tester97):

\[\frac{ x }{ 3 } - \frac{ x-1 }{ 2 }\ge?\]Is this it?

OpenStudy (anonymous):

it is 1

OpenStudy (anonymous):

then you can's solve you can simplify you can rearrange but no solve the best i can offer is this \[\frac{ -x +3}{6 }\]

OpenStudy (anonymous):

that is my bad

OpenStudy (anonymous):

\[x \le 2\]

OpenStudy (anonymous):

no that is wrong